# Robust First and Second-Order Differentiation for Regularized Optimal Transport

### v1

@@ 5 Hessian computation: runtime, accuracy, and success rate
We use the point-cloud datasets sampled from the uniform distribution in unit square again. The task is to calculate the Hessian tensor $\mathcal{T}$ of $\text{OT}_{\epsilon}(C^{\boldsymbol{Y}\rightarrow\boldsymbol{Y}},\boldsymbol{\mu},\boldsymbol{\mu})$ respect to the source data $\boldsymbol{Y}$, where $\boldsymbol{\mu}=\frac{1}{N}\mathbbm{1}_{N}$. By proposition 4, the Hessian satisfies the marginal identity, i.e., $\sum_{k=1}^{M}\mathcal{T}_{k\cdot s\cdot}=2\mu_{s}\mathbb{I}_{d}$. We evaluate the accuracy of the computed Hessian by the marginal error:
@@
Figure 2: Comparison of runtime (in seconds) and marginal error for Hessian computing $\frac{d^{2}\text{OT}_{\epsilon}(C^{\boldsymbol{Y}\rightarrow\boldsymbol{Y}},\boldsymbol{\mu},\boldsymbol{\mu})}{d\boldsymbol{Y}^{2}}$ among three approaches: unroll, implicit differentiation and analytic expression with regularization (ours).
@@ 6 Applications to Shuffled Regression
In this section, we apply our proposed algorithms to solve the shuffled regression problem introduced earlier. It is formulated as the multivariate regression model $\boldsymbol{y}^{*}=\boldsymbol{x}\theta+\xi$, where $\boldsymbol{x}\in\mathbb{R}^{D},\boldsymbol{y}^{*}\in\mathbb{R}^{d},\theta\in\mathbb{R}^{D\times d}$ and $\xi$ is the Gaussian noise independent of $x$. The correspondence between $(\boldsymbol{X},\boldsymbol{Y}^{*})$ is missing. Our goal is to estimate the optimal $\theta^{*}$ using EOT distance as the loss function in the (1). This approach generalizes to unbalanced datasets without requiring $\boldsymbol{X}$ and $\boldsymbol{Y}^{*}$ to have the same number of rows. The gradient and the Hessian of the EOT distance with respect to the parameters $\theta$ are simplified as
$\frac{d\text{OT}_{\epsilon}(C_{\theta},\boldsymbol{\mu},\boldsymbol{\nu})}{d\theta}=\boldsymbol{X}^{\top}\frac{d\text{OT}_{\epsilon}(C_{\theta},\boldsymbol{\mu},\boldsymbol{\nu})}{d\boldsymbol{Y}},$
$\left(\frac{d^{2}\text{OT}_{\epsilon}(C_{\theta},\boldsymbol{\mu},\boldsymbol{\nu})}{d\theta^{2}}\right)_{mtnl}=\sum_{k=1}^{M}\sum_{s=1}^{M}\boldsymbol{X}_{sm}\mathcal{T}_{ktsl}\boldsymbol{X}_{kn}$ (60)
for $t,l=1,\dots,d$ and $m,n=1,\dots,D$. The EOT distance is generally non-convex with respect to $\theta$, so optimization may not converge to the optimal $\theta^{*}$. Our focus is on the convergence speed to a local minimum. First-order methods may converge to a local minimum but require many iterations due to the complicated landscape of the loss function. To accelerate optimization, we propose a two-stage approach. First, we use stochastic gradient descent (SGD) with a random subset of $\boldsymbol{X}$ and the full batch of $\boldsymbol{Y}^{*}$ to quickly approach the local minimum. Then, we switch to a relaxed Newton’s method, using the updated parameter $\hat{\theta}$ as the initial condition. The relaxed-Newton’s method uses step-size $\gamma<1$. In practise, we switch from SGD to relaxed-Newton when the computed Hessian $\frac{d^{2}\text{OT}_{\epsilon}(C_{\theta},\boldsymbol{\mu},\boldsymbol{\nu})}{d\theta^{2}}$ is positive definite. The algorithm is summarized in Algorithm 2.
@@
Input: Data $\boldsymbol{X}$, target data $\boldsymbol{Y}^{*}$, entropy regularization strength $\epsilon$, truncated SVD threshold $\alpha$; initial guess of $\theta^{(0)}$, SGD learning rate $r_{s}$, mini batch size $n_{s}$, maximum epochs $T$; Relaxed Newton learning rate $r_{n}$.
Output: Estimated optimal $\theta^{*}$, regularized optimal transport loss $\text{OT}_{\epsilon}(C_{\theta^{*}},\boldsymbol{\mu},\boldsymbol{\nu})$.
1:Set $\boldsymbol{\tilde{\mu}}\leftarrow\frac{1}{n_{s}}\mathbbm{1}_{n_{s}}$, $\boldsymbol{\mu}\leftarrow\frac{1}{N}\mathbbm{1}_{N}$ and $\boldsymbol{\nu}\leftarrow\frac{1}{N}\mathbbm{1}_{N}$. Stage 1 SGD :
@@
3:     Randomly sample $n_{s}$ rows of $\boldsymbol{X}$, denote as $\boldsymbol{\tilde{X}}$.
4:     Compute $\boldsymbol{\tilde{Y}}\leftarrow\boldsymbol{\tilde{X}}\cdot\theta^{(t)}$ and cost matrix $C^{\boldsymbol{\tilde{Y}}\rightarrow\boldsymbol{Y}^{*}}_{ij}\leftarrow\|\boldsymbol{\tilde{y}}_{i}-\boldsymbol{y}^{*}_{j}\|_{2}^{2}$.
5:     $\theta^{(t+1)}\leftarrow\theta^{(t)}-r_{s}\boldsymbol{\tilde{X}}^{\top}\frac{d\text{OT}_{\epsilon}(C^{\boldsymbol{\tilde{Y}}\rightarrow\boldsymbol{Y}^{*}},\boldsymbol{\tilde{\mu}},\boldsymbol{\nu})}{d\boldsymbol{Y}}$
6:     if  $\frac{d^{2}\text{OT}_{\epsilon}(C_{\theta},\boldsymbol{\mu},\boldsymbol{\nu})}{d\theta^{2}}|_{\theta=\theta^{(t+1)}}$ is positive definite  then
@@
10:     Compute $\boldsymbol{Y}\leftarrow\boldsymbol{X}\cdot\theta^{(t)}$ and cost matrix $C_{ij}\leftarrow\|\boldsymbol{y}_{i}-\boldsymbol{y}^{*}_{j}\|_{2}^{2}$
11:     $\theta^{(t+1)}\leftarrow\theta^{(t)}-r_{n}\left(\frac{d^{2}\text{OT}_{\epsilon}(C_{\theta^{(t)}},\boldsymbol{\mu},\boldsymbol{\nu})}{d\theta^{2}}\right)^{-1}\left(\boldsymbol{\tilde{X}}^{\top}\frac{d\text{OT}_{\epsilon}(C_{\theta^{(t)}},\boldsymbol{\mu},\boldsymbol{\nu})}{d\boldsymbol{Y}}\right)$
12:     if $\text{OT}_{\epsilon}(C_{\theta^{(t+1)}},\boldsymbol{\mu},\boldsymbol{\nu})$ doesn’t improve then
@@ 6.1 Shuffled Regression with Gaussian Mixtures
We first generate $N=500$ data points $\boldsymbol{X}\in\mathbb{R}^{5}$ from a Gaussian mixture distribution with three clusters whose standard deviation is $[0.3,0.05,0.6]$. The parameter $\theta^{*}\in\mathbb{R}^{5\times 2}$ is generated with components $\theta^{*}_{mt}\sim\mathcal{N}(0,1)$, and the Gaussian noise $\boldsymbol{\xi}\in\mathbb{R}^{2}$ follows $\boldsymbol{\xi}\sim\mathcal{N}(\mathbb{0},0.04\mathbb{I}_{2})$. We then compute $\boldsymbol{y}_{i}^{*}=\boldsymbol{x}_{i}\theta^{*}+\boldsymbol{\xi}_{i}$, randomly and completely permute the order of $\boldsymbol{y}_{i}^{*}$, removing $\boldsymbol{X}$-to-$\boldsymbol{Y}^{*}$ correspondence.
Starting with an random initial condition $\theta^{(0)}$ from the standard normal distribution, the target data $\boldsymbol{Y}^{*}$ and the initial data $\boldsymbol{Y}(\theta^{(0)})$ are shown in Figure 3(a). We use the two-stage algorithm described in Algorithm 2. In the first stage, we perform 10 iterations of SGD on 100 random source data points with a learning rate of 0.001. In the second stage, we use a relaxed Newton’s method with a learning rate of 0.5. We compare this to a gradient descent (GD) method with a learning rate of 0.001.
Figure 3(b) shows that both methods correctly map the data $\boldsymbol{X}$ to the target data $\boldsymbol{Y}^{*}$. Figure 3(c-d) shows that both methods converge to the optimal $\theta^{*}$, but the relaxed Newton’s method is faster and more accurate. The relaxed Newton’s method converges in 12 iterations with a runtime of 2.35 seconds, while GD takes 2000 iterations and 64.77 seconds, which is 27 times longer. Additionally, the relaxed Newton’s method achieves nearly one order of magnitude better accuracy in terms of the $L_{2}$ error $\|\theta-\theta^{*}\|_{2}$.
@@
Using the MobilNet10 dataset [25], we create a study room with a chair (500 points), a desk (1500 points), and a sofa (1500 points), denoted as $\boldsymbol{X}$. We apply a linear transformation including random rotation and scaling, and add Gaussian noise: $\boldsymbol{Y}^{*}=\boldsymbol{X}\theta^{*}+\boldsymbol{\xi}$ with $\boldsymbol{\xi}\sim\mathcal{N}(0,4\times 10^{-4}\mathbb{I}_{3})$. The rows of $\boldsymbol{Y}^{*}$ are randomly permuted to remove correspondence.
We use the algorithm from Algorithm 2. The initial parameter $\theta^{(0)}$ is a standard Gaussian perturbation of the optimal parameter $\theta^{*}$. In the first stage, we perform 5 iterations of SGD on 500 random data points with a learning rate of 0.1. In the second stage, we use a relaxed Newton’s method with a learning rate of 0.5. For comparison, the GD method uses a learning rate of 0.1.
@@ 7 Conclusion
In this work, we computed first-order and second-order derivatives for the parameterized regularized optimal transport (OT) distance. Specifically, we derived explicit analytical expressions for the gradient of the Sinkhorn distance and the Hessian of the entropy-regularized OT (EOT) distance with respect to the source data $\boldsymbol{Y}$.
To address the numerical instability and high memory consumption typically associated with Hessian computation in large-scale, multi-dimensional problems, we developed a fast, stable, and memory-efficient algorithm using spectral analysis of the ill-posed linear system. Our algorithm demonstrated significant improvements in both efficiency and accuracy on various benchmark datasets.
These results highlight the potential of our proposed algorithm to enhance the performance and reliability of optimization tasks in complex, high-dimensional spaces, particularly in regression without correspondence.
Future work may explore further refinements of our stabilization thresholds by studying the limiting behavior of condition numbers on other random datasets, as well as the computation of robust second-order differentiation for more general regularized and constrained optimal transport problems.
@@
The software package implementing the proposed algorithms can be found on:
https://github.com/yexf308/OTT-Hessian.
@@ Appendix A Proofs for spectral analysis
The proof of Proposition 11 is based on an application of Weyl’s inequality to study the eigenvalues of the $H$-matrix under perturbation.

### v2

@@ 5 Hessian computation: runtime, accuracy, and success rate
We use the point-cloud datasets sampled from the uniform distribution in unit square again. The task is to calculate the Hessian tensor $\mathcal{T}$ of $\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\mu})$ respect to the source data $\mathbf{Y}$, where $\bm{\mu}=\frac{1}{N}\mathbb{1}_{N}$. By proposition 5, the Hessian satisfies the marginal identity, i.e., $\sum_{k=1}^{M}\mathcal{T}_{k\cdot s\cdot}=2\mu_{s}\mathbb{I}_{d}$. We evaluate the accuracy of the computed Hessian by the marginal error:
@@
Figure 2: Comparison of runtime (in seconds) and marginal error for Hessian computing $\frac{d^{2}\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\mu})}{d\mathbf{Y}^{2}}$ among three approaches: unroll, implicit differentiation and analytic expression with regularization (ours).
@@ 6 Applications to Shuffled Regression
In this section, we apply our proposed algorithms to solve shuffled regression problems. It is formulated as the multivariate regression model $\mathbf{y}^{*}=\mathbf{x}\theta+\bm{\xi}$, where $\mathbf{x}\in\mathbb{R}^{D},\mathbf{y}^{*}\in\mathbb{R}^{d},\theta\in\mathbb{R}^{D\times d}$ and $\bm{\xi}$ is the Gaussian noise independent of $\mathbf{x}$. The correspondence between $(\mathbf{X},\mathbf{Y}^{*})$ is missing. Our goal is to estimate the optimal $\theta^{*}$ using EOT distance as the loss function in the (1). We choose both probability weights $\bm{\mu}$ and $\bm{\nu}$ as uniform weights. This approach generalizes to datasets without requiring $\mathbf{X}$ and $\mathbf{Y}^{*}$ to have the same number of rows. The gradient and the Hessian of the EOT distance with respect to the parameters $\theta$ are simplified as
$\frac{d\text{OT}_{\epsilon}(\mathbf{C}_{\theta},\bm{\mu},\bm{\nu})}{d\theta}=\mathbf{X}^{\top}\frac{d\text{OT}_{\epsilon}(\mathbf{C}_{\theta},\bm{\mu},\bm{\nu})}{d\mathbf{Y}},$
$\left(\frac{d^{2}\text{OT}_{\epsilon}(\mathbf{C}_{\theta},\bm{\mu},\bm{\nu})}{d\theta^{2}}\right)_{mtnl}=\sum_{k=1}^{M}\sum_{s=1}^{M}\mathbf{X}_{sm}\mathcal{T}_{ktsl}\mathbf{X}_{kn}$ (54)
for $t,l=1,\dots,d$ and $m,n=1,\dots,D$. The EOT distance is generally non-convex with respect to $\theta$, so optimization may not converge to the optimal $\theta^{*}$. Our focus is on the convergence speed to a local minimum. First-order methods may converge to a local minimum but require many iterations due to the complicated landscape of the loss function. To accelerate optimization, we propose a two-stage approach. First, we use stochastic gradient descent (SGD) with a random subset of $\mathbf{X}$ and the full batch of $\mathbf{Y}^{*}$ to quickly approach the local minimum. Then, we switch to a relaxed Newton’s method, using the updated parameter $\hat{\theta}$ as the initial condition. The relaxed-Newton’s method uses step-size $\gamma<1$. In practise, we switch from SGD to relaxed-Newton when the computed Hessian $\frac{d^{2}\text{OT}_{\epsilon}(\mathbf{C}_{\theta},\bm{\mu},\bm{\nu})}{d\theta^{2}}$ is positive definite. The algorithm is summarized in Algorithm 2. In the situation that the Hessian is too expansive to calculate multiple times, we can use the Hessian calculated from symmetric entropic OT problem of $\mathbf{Y}^{*}$ to $\mathbf{Y}^{*}$, as the approximation of the true Hessian in Newton’s method.
@@
Input: Data $\bm{X}$, target data $\bm{Y}^{*}$, entropy regularization strength $\epsilon$, truncated SVD threshold $\alpha$; initial guess of $\theta^{(0)}$, SGD learning rate $r_{s}$, mini batch size $n_{s}$, maximum epochs $T$; Relaxed Newton learning rate $r_{n}$.
Output: Estimated optimal $\theta^{*}$, regularized optimal transport loss $\text{OT}_{\epsilon}(C_{\theta^{*}},\bm{\mu},\bm{\nu})$.
1:Set $\bm{\tilde{\mu}}\leftarrow\frac{1}{n_{s}}\mathbb{1}_{n_{s}}$, $\bm{\mu}\leftarrow\frac{1}{N}\mathbb{1}_{N}$ and $\bm{\nu}\leftarrow\frac{1}{N}\mathbb{1}_{N}$. Stage 1 SGD :
@@
3:     Randomly sample $n_{s}$ rows of $\bm{X}$, denote as $\bm{\tilde{X}}$.
4:     Compute $\bm{\tilde{Y}}\leftarrow\bm{\tilde{X}}\cdot\theta^{(t)}$ and cost matrix $C^{\bm{\tilde{Y}}\rightarrow\bm{Y}^{*}}_{ij}\leftarrow\|\bm{\tilde{y}}_{i}-\bm{y}^{*}_{j}\|_{2}^{2}$.
5:     $\theta^{(t+1)}\leftarrow\theta^{(t)}-r_{s}\bm{\tilde{X}}^{\top}\frac{d\text{OT}_{\epsilon}(C^{\bm{\tilde{Y}}\rightarrow\bm{Y}^{*}},\bm{\tilde{\mu}},\bm{\nu})}{d\bm{Y}}$
6:     if  $\frac{d^{2}\text{OT}_{\epsilon}(C_{\theta},\bm{\mu},\bm{\nu})}{d\theta^{2}}|_{\theta=\theta^{(t+1)}}$ is positive definite  then
@@
10:     Compute $\bm{Y}\leftarrow\bm{X}\cdot\theta^{(t)}$ and cost matrix $C_{ij}\leftarrow\|\bm{y}_{i}-\bm{y}^{*}_{j}\|_{2}^{2}$
11:     $\theta^{(t+1)}\leftarrow\theta^{(t)}-r_{n}\left(\frac{d^{2}\text{OT}_{\epsilon}(C_{\theta^{(t)}},\bm{\mu},\bm{\nu})}{d\theta^{2}}\right)^{-1}\left(\bm{\tilde{X}}^{\top}\frac{d\text{OT}_{\epsilon}(C_{\theta^{(t)}},\bm{\mu},\bm{\nu})}{d\bm{Y}}\right)$
12:     if $\text{OT}_{\epsilon}(C_{\theta^{(t+1)}},\bm{\mu},\bm{\nu})$ doesn’t improve then
@@ 6.1 Shuffled Regression with Gaussian Mixtures
We first generate $N=500$ data points $\mathbf{X}\in\mathbb{R}^{5}$ from a Gaussian mixture distribution with three clusters. The mean vectors for these clusters are randomly chosen, and the covariance matrices are $0.3^{2}\mathbb{I}_{5},0.05^{2}\mathbb{I}_{5}$ and $0.6^{2}\mathbb{I}_{5}$, respectively. The parameter $\theta^{*}\in\mathbb{R}^{5\times 2}$ is generated with components $\theta^{*}_{mt}\sim\mathcal{N}(0,1)$, and the Gaussian noise $\bm{\xi}\in\mathbb{R}^{2}$ follows $\bm{\xi}\sim\mathcal{N}(\mathbb{0},0.04\mathbb{I}_{2})$. We then compute $\mathbf{y}_{i}^{*}=\mathbf{x}_{i}\theta^{*}+\bm{\xi}_{i}$, randomly and completely permute the order of $\mathbf{y}_{i}^{*}$, removing $\mathbf{X}$-to-$\mathbf{Y}^{*}$ correspondence.
Starting with an random initial condition $\theta^{(0)}$ from the standard normal distribution, the target data $\mathbf{Y}^{*}$ and the initial data $\mathbf{Y}_{\theta^{(0)}}$ are shown in Figure 3(a). We use the two-stage algorithm described above. In the first stage, we perform 10 iterations of SGD on 100 random source data points with a learning rate of 0.001. In the second stage, we use a relaxed Newton’s method with a learning rate of 0.5. We compare this to a gradient descent (GD) method with a learning rate of 0.001.
Figure 3(b) shows that both methods correctly map the data $\mathbf{X}$ to the target data $\mathbf{Y}^{*}$. Figure 3(c-d) shows that both methods converge to the optimal $\theta^{*}$, but the relaxed Newton’s method is faster and more accurate. The relaxed Newton’s method converges in 12 iterations with a runtime of 2.35 seconds, while GD takes 2000 iterations and 64.77 seconds, which is 27 times longer. Additionally, the relaxed Newton’s method achieves nearly one order of magnitude better accuracy in terms of the $L_{2}$ error $\|\theta-\theta^{*}\|_{2}$.
@@
Using the MobilNet10 dataset [27], we create a study room with a chair (500 points), a desk (1500 points), and a sofa (1500 points), denoted as $\mathbf{X}$. We apply a linear transformation including random rotation and scaling, and add Gaussian noise: $\mathbf{Y}^{*}=\mathbf{X}\theta^{*}+\bm{\xi}$ with $\bm{\xi}\sim\mathcal{N}(0,4\times 10^{-4}\mathbb{I}_{3})$. The rows of $\mathbf{Y}^{*}$ are randomly permuted to remove correspondence.
We use the two-stage algorithm described above. The initial parameter $\theta^{(0)}$ is a standard Gaussian perturbation of the optimal parameter $\theta^{*}$. In the first stage, we perform 5 iterations of SGD on 500 random data points with a learning rate of 0.1. In the second stage, we use a relaxed Newton’s method with a learning rate of 0.5. For comparison, the GD method uses a learning rate of 0.1.
@@ 7 Conclusion
In this work, we computed first-order and second-order derivatives for the parameterized regularized optimal transport (OT) distance. Specifically, we derived explicit analytical expressions for the gradient of the Sinkhorn distance and the Hessian of the entropy-regularized OT (EOT) distance with respect to the source data $\mathbf{Y}$.
To address the numerical instability and high memory consumption typically associated with Hessian computation in large-scale, multi-dimensional problems, we developed a fast, stable, and memory-efficient algorithm using spectral analysis of the ill-posed linear system. Our algorithm demonstrated significant improvements in both efficiency and accuracy on benchmark datasets.
Future work may explore further accelerating the computation by sparse approximation of the Hessian in the computation of both the OT distance and its second-order derivatives, spectral analysis of the H-matrix with random datasets with a focus on the asymptotic behavior in the large $N$ or small $\epsilon$ limit, and robust second-order differentiation for general regularized and constrained optimal transport problems.
@@
The software package implementing the proposed algorithms can be found on https://github.com/yexf308/OTT-Hessian.
@@ Appendix A Proofs for Spectral Analysis
A.1 Proof of Lemma 7
Proof.
First, we observe that
$\text{diag}(\bm{\nu})^{-1}\bm{\Pi}^{\top}\text{diag}(\bm{\mu})^{-1}\bm{\Pi}\mathbb{1}_{N}=\text{diag}(\bm{\nu})^{-1}\bm{\Pi}^{\top}\text{diag}(\bm{\mu})^{-1}\bm{\mu}=\text{diag}(\bm{\nu})^{-1}\bm{\Pi}^{\top}\mathbb{1}_{M}=\mathbb{1}_{N}$
This shows the eigenvector of $\lambda=1$ is $\mathbb{1}_{N}$.
Furthermore, we show that $\lambda=1$ has multiplicity one by showing that any solution $\bm{v}$ to $\bm{v}=\text{diag}(\bm{\nu})^{-1}\bm{\Pi}^{\top}\text{diag}(\bm{\mu})^{-1}\bm{\Pi}\bm{v}$, must have identical entries. Suppose that there exists a $\bm{v}=(v_{1},\ldots,v_{N})$ has non-constant entries. Let $\widetilde{\bm{v}}$ denote $\widetilde{\bm{v}}=\text{diag}(\bm{\mu})^{-1}\bm{\Pi}{\bm{v}}$. Then, we have $\max_{1\leq i\leq M}\widetilde{v}_{i}<\max_{1\leq j\leq N}v_{j}$, because the entries of $\bm{\Pi}$ are all positive, $\bm{\Pi}\mathbb{1}_{N}=\mu$ and each entry of $\text{diag}(\bm{\mu})^{-1}\bm{\Pi}{\bm{v}}$ is a weighted average of $v$. Similarly, we have $\max_{1\leq j\leq N}[\text{diag}(\bm{\nu})^{-1}\bm{\Pi}^{\top}\widetilde{\bm{v}}]_{j}<\max_{1\leq i\leq M}\widetilde{v}_{i}$. Since $\bm{v}=\text{diag}(\bm{\nu})^{-1}\bm{\Pi}^{\top}\widetilde{\bm{v}}$ , we obtain a contradiction: $\max_{1\leq j\leq N}v_{j}<\max_{1\leq j\leq N}v_{j}$. Hence, $\bm{v}=(v_{1},\ldots,v_{N})$ must have constant entries, the sames as $\mathbb{1}_{N}$ up to a scalar factor.
Next, we show that all eigenvalues must be no greater than one. Suppose that $\lambda$ is an eigenvalue with $v$ being its normalized eigenvector. Thus, denoting $u_{1}=\bm{\Pi}\text{diag}(\bm{\nu})^{-1}v$ and $u_{2}=\text{diag}(\bm{\mu})^{-1}\bm{\Pi}{\bm{v}}$, we have
$\lambda=\frac{{\bm{v}}^{\top}\text{diag}(\bm{\nu})^{-1}\bm{\Pi}^{\top}\text{diag}(\bm{\mu})^{-1}\bm{\Pi}{\bm{v}}}{{\bm{v}}^{\top}\bm{v}}=\langle{u_{1},u_{2}}\rangle\leq\|u_{1}\|\|u_{2}\|\leq 1,$
where the last inequality follows from the facts that $\|u_{1}\|^{2}=\|\bm{\Pi}\text{diag}(\bm{\nu})^{-1}\bm{v}\|_{2}^{2}\leq 1$, and that $\|u_{2}\|=\|\text{diag}(\bm{\mu})^{-1}\bm{\Pi}{\bm{v}}\|\leq 1$.
A similar argument shows that $\text{diag}(\bm{\mu})^{-1}\bm{\Pi}\text{diag}(\bm{\nu})^{-1}\bm{\Pi}^{\top}$ must have simple eigenvalue $\lambda=1$ and eigenvector $\mathbb{1}_{M}$. ∎
A.2 Proof of Proposition 12
The proof of Proposition 12 is based on an application of Weyl’s inequality to study the eigenvalues of the $\mathbf{H}$-matrix under perturbation. The result is of general interest beyond this study, so we state it as a lemma.